If Sylvia works overtime, she is paid 1 1/2 times as much per hour as usual. After working her usual 40 h last week, she worked an additional 4 h over-time. If she made $552 last week, find her usual hourly wage.
Answer by nyc_function(2741)

Let h = Sylvia's hourly wage
Let 1_1/2 = 3/2
40x + 4(3/2) = 552
Solve equation for h and you will find Sylvia's hourly wage.